# Comparison of marker-less 2D image-based methods for infant pose estimation

### v2

@@ Abstract
In this study we compare the performance of available generic- and infant-pose estimators for a video-based automated general movement assessment (GMA), and the choice of viewing angle for optimal recordings, i.e., conventional diagonal view used in GMA vs. top-down view. We used 4500 annotated video-frames from 75 recordings of infant spontaneous motor functions from 4 to 26 weeks. To determine which pose estimation method and camera angle yield the best pose estimation accuracy on infants in a GMA related setting, the distance to human annotations and the percentage of correct key-points (PCK) were computed and compared. The results show that the best performing generic model trained on adults, ViTPose, also performs best on infants. We see no improvement from using infant-pose estimators over the generic pose estimators on our infant dataset. However, when retraining a generic model on our data, there is a significant improvement in pose estimation accuracy. The pose estimation accuracy obtained from the top-down view is significantly better than that obtained from the diagonal view, especially for the detection of the hip key-points. The results also indicate limited generalization capabilities of infant-pose estimators to other infant datasets, which hints that one should be careful when choosing infant pose estimators and using them on infant datasets which they were not trained on. While the standard GMA method uses a diagonal view for assessment, pose estimation accuracy significantly improves using a top-down view. This suggests that a top-down view should be included in recording setups for automated GMA research.
{IEEEkeywords}
deep neural networks, full body pose estimation, GMA, infant motion analysis
@@ 1 Introduction
When classic biomarker approaches fail to detect developmental conditions early or predict neurodevelopmental outcomes following pre- or perinatal brain lesions or complications during pregnancy, assessments of overt neurofunctions come into play [1, 2, 3]. An early detection of biological markers of adverse outcomes or early diagnosis of developmental conditions facilitates early intervention, when brain plasticity is at its peak, aiming to achieve best possible long-term outcomes [4]. One of the available tools to functionally assess the developing nervous system is the Prechtl general movements assessment (GMA) [5]. Over the last decades, GMA widened its scope and applicability to serve as a general estimate of the integrity of the developing nervous system. It became a tool to look beyond the mere distinction of typical development vs. high risk for cerebral palsy, e.g., to determine the impact of viral infections on postnatal neurodevelopment or describe the early phases of neurodevelopmental conditions such as autism [6, 7]. This broad applicability was the reason we used the multi-camera recording technique which included the standard GMA perspective as a use case for this study [8, 7, 9]. GMA is a gestalt-based observational tool to classify spontaneous infant motor functions in the first months of life [5]. Infants up to 20 weeks of post-term age are positioned in prone in a cot and their spontaneous movements, i.e. unstimulated, are video-recorded. These movements can be analyzed and classified into physiological or atypical movement patterns [5, 10, 2, 3]. Given its high sensitivity and specificity it has become one tool of choice for the early detection of cerebral palsy in the postnatal period [3]. Recently, efforts have been intensified to automate this clinical method using computers (for recent reviews see [11, 12, 13, 14, 15, 16]).
In early studies, spontaneous movements in the first months of life were measured with sensors directly attached to the infant [17, 18], or using optical flow methods [19, 20, 21]. With the advent of neural networks for pose estimation [22, 23, 24, 25] current studies have switched to using skeleton keypoints for visual-based movement analysis and classification [26, 27, 28, 29, 30, 31, 32, 33]. At first, those pose estimators were only trained on adults and used for infants without modification. Later work identified problems with this approach and developed infant specific models [34, 35, 36, 37].
@@
Overall, it is difficult to compare existing models, evaluate their true performance and generalize their results. The main reason is, that models were trained and tested on silo-data-sets (patient data) as we are facing a complex data sharing issue which presents a true obstacle for all video-based clinical methods [30]. There are comparisons between related model architectures when selecting the best model [35, 36], but there has not yet been any comparison of completely different models that were trained on completely different training sets. Thus, it is unknown if the generalization capabilities of special infant pose estimators are sufficient for the use on datasets the particular models were not trained on. To test the generalization capabilities of generic pose estimation models, and to compare those to the infant specific models, we first analyzed four different pose estimation models, OpenPose [38, 39, 23], MediaPipe [40, 41], HRNet [42, 43] and ViTPose [25]. In a second step, we retrained the best performing model ViTPose [44] on our dataset. We then compare the retrained model to the two infant pose estimators, AggPose [35] and AGMA-HRNet48 [36].
Furthermore, most analysis so far has been carried out in 2D space and has only recently been transferred to the 3D space [45, 36, 46, 37]. We have also collected a new multi-view infant video dataset, taken from four different camera angles (see Figure 1a) which allows 3D reconstruction. We chose these four angles to provide occlusion robustness for various infant poses. The viewing angle has great influence on the visibility of different body parts and should therefore be carefully considered when designing an experimental setup. Since we captured four angles at once, this offers the opportunity to directly compare the different viewing angles that were commonly used in other works, a diagonal view or a top view (see Figure 1b,c), and quantify their influence on pose estimation accuracy. GMA, in its clinical application, is usually done with single 2D-RGB cameras in a diagonal view [8, 47]. For this study, we split the analyses done on generic and specific infant pose estimators by viewing angle and compare them. We also specifically retrained on single angles trying to achieve the best performing models.
The results can be used as an indication for the optimal selection of camera angles in future projects.
The main contributions of this work are the comparison of state-of-the-art infant pose estimators and the influence of the viewing angle on pose estimation accuracy. Specifically, we address four main research questions:
  1. 1.
@@
  2. 2.
@@
  3. 3.
@@
  4. 4.
@@ 2.1.1 Participants
Participants included in this study were prospectively recruited from 2021 to 2023 in Göttingen, Germany, and its close surroundings. The umbrella project aims to investigate cross-domain ontogenetic development in early infancy. To embrace the variability of the targeted dimension, i.e., spontaneous motor functions, 31 participants (17 female) were included. The gestational age at birth of the sample ranged from 34 to 42 weeks. At the time of data analysis, no participant was diagnosed with a neurological or neuromotor deficit, nor any neurodevelopmental impairment. All parents of the participating infants provided written informed consent to study participation and publication of depersonalized data and results. The study was approved by the Institutional Review Board (Ethics Commission) of the University Medical Center Göttingen (19/2019).
@@ 2.1.2 Infant movements recordings
Standard laboratory recordings of infant movements at three timepoints (T1 – T3) were included in the current study. These were extracted from data of the umbrella project, assessments, among others, infants’ spontaneous movements in a standard laboratory-setting from the 4th to the 18th week of post-term age (PTA; corrected age; from here on, ages refer to the post-term age if not otherwise specified). For the current study, available laboratory recordings (n = 75) of the participants for the following timepoints are analyzed: T1: 28 ± 2 days, T2: 84 ± 2 days, and T3: 112 ± 2 days of PTA. As known, infants at 4 weeks (corresponding to T1) present a different spontaneous movement pattern than at 12 and 16 weeks (corresponding to T2 and T3; [49, 50]). With standard recordings from T1 to T3, we intended to cover the distinct age-specific spontaneous movement repertoires of the young infants [10].
@@ 2.1.4 Image selection and annotation
The 75 recordings were split into two parts of 50 and 25 which were annotated by one research assistant each, using JARVIS Annotation Tool [51]. Additionally, 10% of the dataset, regardless of the split, were double-annotated for human error and individual joint labeling difficulty estimation.
@@ 2.2.1 Generic pose estimation
was one of the first pose estimation frameworks used for extraction of movements for analysis and classification of infant motor functions [26, 52, 27]. We included this framework as a baseline method to evaluate the performance gain of state-of-the-art models.
@@
This framework is not primarily intended for accurate pose estimation, but rather for fast inference on mobile devices. It was interesting for our comparison as Google advertises it as suitable for Yoga and fitness applications and included 25000 images of fitness exercises, so we expected it to deal well with different viewing angles and poses.
@@
The HR in HRNet stands for high resolution. The particular version used has an input size of $512\times 512$ pixels. Because our videos are of sufficient resolution and adequate quality with low compression artefacts, this could benefit the pose estimation accuracy.
@@
This model is the current best performing framework for human pose estimation on the COCO test-dev dataset and uses a Vision Transformer architecture instead of the conventional CNN.
@@ 2.2.2 Infant pose estimation
The AggPose model is based on the Transformer architecture, like ViTPose, and is specifically trained for infant pose detection on a proprietary dataset. In general, the images in the dataset also stem from recordings used for GMA, with infants in supine position, like in our dataset.
@@
Soualmi et al. also retrained infant pose estimation networks on a proprietary dataset of GMA recordings. Compared to our dataset, the images contain clinical equipment (e.g., tubes, electrodes, cables), leading to more visual clutter in the scenes. The authors retrained multiple networks, of which we chose the HRNet48, because it is the same architecture as the one we used for the comparison of the generic pose estimation frameworks.
@@ 2.2.3 Retraining
We retrained the best performing framework ViTPose (see Section 3.2) on our labeled dataset using 5-fold cross-validation. We manually split the dataset such that all images of each infant were contained in one fold only. This ensures the networks do not get better test results by overfitting on specific infants, which are correlated across different recording dates. Each of the five splits therefore contained 900 frames, made up from 450 different situations recorded from two different angles. Because of the different numbers of recordings done for each infant, four splits contained images from six different infants and one split contained images from seven different infants. For training, we randomly split 10% of the training set off as validation set, yielding a final training set size of 3240 frames in each fold. For each fold, We then retrained the model for a maximum of 50 epochs, keeping the model with the highest PCK on the validation set for evaluation on the test set.
2.3 Quantification measures
It is not possible to transform any of the quantification measures directly into real-world distances without using 3D triangulation, which was not done in this work, or assuming the points are on the same plane, which is untrue because the infants frequently lift their extremities. However, one pixel on the plane of the bed, always corresponds to a distance $\leq 0.8$ mm. Therefore, 80% of the difference to annotation is an upper bound to the real-world distance in mm. Considering the mean torso length of 306 pixels, PCK@0.1 yields the percentage of points that were detected within approximately 2.5cm of the ground truth (given 0.8 mm as the upper bound).
@@ 2.3 Quantification measures
2.3.1 Pose estimation error
We used two main evaluation metrics to assess the performance of the pose estimation methods: the difference between predicted keypoint position and the human annotation, $d_{a}$, in pixels, and the percentage of correct keypoints (PCK). The difference between model prediction and the human annotation ($d_{a}$) is the Euclidean distance between the predicted and human label in pixels:
$d_{a}=\sqrt{(p_{x}-h_{x})^{2}+(p_{y}-h_{y})^{2}},$ (1)
where $p_{x/y}$ are the network predictions and $h_{x/y}$ the human labels for a keypoint. For PCK [53], all keypoint predictions are classified as either correct or incorrect, based on their distance to the label (as given above). The PCK score is then defined as the fraction of correct keypoints out of all keypoints. The distance threshold to be considered correct is defined as a percentage of the apparent torso length of the infant in the frame, measured as the distance between the left shoulder and hip. In this work we considered the PCK at 5%, 7.5% and 10% of the torso size, denoted by PCK@0.05, PCK@0.75 and PCK@0.1, respectively.
Our dataset is not biased to any specific side (left or right) of the infants. Moreover, labeling left and right sides of hands, knees, etc. is not inherently different. Therefore, we merged the errors for keypoints of the left and right side into one.

### v3

@@ Abstract
In this study we compare the performance of available generic- and specialized infant-pose estimators for a video-based automated general movement assessment (GMA), and the choice of viewing angle for optimal recordings, i.e., conventional diagonal view used in GMA vs. top-down view. We used 4500 annotated video-frames from 75 recordings of infant spontaneous motor functions from 4 to 16 weeks. To determine which pose estimation method and camera angle yield the best pose estimation accuracy on infants in a GMA related setting, the distance to human annotations and the percentage of correct key-points (PCK) were computed and compared. The results show that the best performing generic model trained on adults, ViTPose, also performs best on infants. We see no improvement from using specific infant-pose estimators over the generic pose estimators on our infant dataset. However, when retraining a generic model on our data, there is a significant improvement in pose estimation accuracy. This indicates limited generalization capabilities of infant-pose estimators to other infant datasets, meaning that one should be careful when choosing infant pose estimators and using them on infant datasets which they were not trained on. The pose estimation accuracy obtained from the top-down view is significantly better than that obtained from the diagonal view (the standard view for GMA). This suggests that a top-down view should be included in recording setups for automated GMA research.
keywords:
full body pose estimation, infant motion analysis, deep neural networks, GMA
@@ 1.1 Background
When classic biomarker approaches fail to detect developmental conditions early or predict neurodevelopmental outcomes following pre- or perinatal brain lesions or complications during pregnancy, assessments of overt neurofunctions come into play [1, 2, 3]. An early detection of biological markers of adverse outcomes or early diagnosis of developmental conditions facilitates early intervention, when brain plasticity is at its peak, aiming to achieve best possible long-term outcomes [4]. The Prechtl general movement assessment (GMA) [5] relies on human visual gestalt perception to delineate the spontaneous motor repertoire which is continuously present from fetal life to about 5 months post-term age [6, 7]. The fetus and the newborn move without sensory stimulation in age-specific patterns, so called general movements (GMs). These movements involve the whole body, appear as writhing movements from about 6-9 weeks post-term age and change their character to so-called fidgety movements later on. Neurobiology has coined the term ‘central pattern generator’ (CPG) for the underlying neural circuitry of endogenously generated activity that is not triggered by sensory stimulation but can be modulated by the periphery. General movements that present with a variable sequence of arm, leg, neck, and trunk movements have been studied through human and augmented Gestalt perception approaches [8, 9, 10, 11]. Over the last decades, GMA widened its scope and applicability to serve as a general estimate of the integrity of the developing nervous system. It became a tool to look beyond the mere distinction of typical development vs. high risk for cerebral palsy, e.g., to determine the impact of viral infections on postnatal neurodevelopment or describe the early phases of neurodevelopmental conditions such as autism [12, 13]. This broad applicability was the reason we used the multi-camera recording technique for a population based study which included the standard GMA perspective as a use case [14, 13, 10]. GMA is a gestalt-based observational tool to classify spontaneous infant motor functions in the first months of life [5]. Infants up to 20 weeks of post-term age are positioned in supine position in a cot and their spontaneous movements, i.e. unstimulated, are video-recorded. These movements can be analyzed and classified into physiological or atypical movement patterns [5, 6, 2, 3]. Given its high sensitivity and specificity GMA has become one tool of choice for the early detection of cerebral palsy in the early postnatal period [3].
1.2 State of the Art
Recently, efforts have been intensified to automate GMA using computers (for recent reviews see [15, 9, 16, 17, 18, 19]). In early studies, spontaneous movements in the first months of life were measured with sensors directly attached to the infant [20, 21]. However, the required procedures and the presence of the sensors could potentially change motion [9], thus, efforts to augment or replace wearable sensors with non-invasive computer vision based methods have become increasingly popular. Early camera based works used optical flow methods [22, 23, 24, 25] for motion detection. With the advent of neural networks for pose estimation [26, 27, 28, 29] current studies have switched to pose estimation and skeleton keypoints for visual-based movement analysis and classification [30, 31, 32, 33, 34, 35, 36, 37, 38]. At first, most automated GMA studies [30, 31, 32, 35] used OpenPose [45, 46, 27]. More recent approaches changed to newer pose estimators like EfficientPose [33], the adaptable pose estimation framework DeepLabCut [34], or a derivative version of OpenPose fine tuned on infant datasets [37, 38].
@@ 1.2 State of the Art
Recent findings revealed ViTPose to be the best performing framework for human pose estimation using the COCO dataset [29]. However, all current pose estimators with large underlying datasets are trained on adults and then used for infant pose estimation without modification. One of the reasons is that there are no large public infant datasets available. The field of infant pose estimation for clinical applications in general faces a complex data sharing issue which presents a true obstacle for all video-based clinical methods [35]. Infant specific models [39, 40, 41, 42, 43] are consequently mostly trained and tested on silo-datasets (patient data) and results have not yet been compared on a common test-set. There are comparisons between related model architectures when selecting the best model [41, 42], but they are all trained on the same data. There is also a comparison of different existing models on infant data[44], however, there, no infant data was used in training. Thus, it is unknown if the generalization capabilities of specialized infant pose estimators are sufficient for use on datasets the particular models were not trained on.
Additionally, GMA, in its clinical application, is usually done with single 2D-RGB cameras in a diagonal view [14, 53]. This viewing angle is uncommon for pictures of adults on which generic pose estimators are trained, which might also have effects on the pose estimators’ performance on infant datasets.
1.3 Contribution
In this work, we compare various infant pose estimators (generic and infant specific) to evaluate the performance increase using state of the art methods. We also analyze the generalization capabilities of infant specific pose estimators on datasets different from training, and quantify the effect of different viewing angles on pose estimation accuracy. This is important, as selection of the best pose estimator will directly influence the quality of any task done on the extracted poses.
Specifically, we address the following four main research questions:
@@ 1.3 Contribution
    To test the generalization capabilities of generic pose estimation models, and to later compare those to the infant specific models, we first analyzed four different pose estimation models, OpenPose [45, 46, 27], MediaPipe [47, 48], HRNet [49, 50] and ViTPose [29]. Specific reasons for the inclusion of all models are given in section 2.2. We find that ViTPose performs best.
@@
    We evaluated the infant-specific pose estimators AggPose [41] and AGMA-HRNet48 [42] on our dataset. In a second step, we retrained the best performing generic model ViTPose [29] on our dataset and then compared the retrained model to the two specific infant pose estimators. As detailed in 1.2, a comparison like this has not yet been done for infants using data from a source different from the one used in training. We find that retraining, as predicted, increases accuracy, but also that the infant-specific models do not generalize well and only perform at or below the level of the best generic model.
@@
    We studied the influence of the viewing angle on the pose estimation methods by splitting the analysis by viewing angles. Since the data used in this study comes from a recording setup with multiple angles (see Figure 1a) for later 3D-reconstruction, we can directly compare the different viewing angles that were commonly used in other works, a diagonal view and a top view (see Figure 1b,c), and quantify their influence on pose estimation accuracy. We find better performance for the top view as compared to the diagonal (GMA preferred) view.
@@
    We specifically retrained ViTPose on single view angles, to test if this further increases the accuracy, or at least leads to equal performance on top and diagonal views (for single model, the diagonal view was worse in any situation). We find that angle specific ViTPose models lead to no improvement over a unified model.
Our findings provide useful guidelines for the optimal selection of camera angles and pose estimation models for future projects on automated infant movement classification (such as GMA).
@@ 2.1.1 Participants
Participants included in this study were prospectively recruited from 2021 to 2023 in Göttingen, Germany, and its close surroundings. The umbrella project aims to investigate cross-domain ontogenetic development in early infancy. To embrace the variability of the targeted dimension, i.e., spontaneous motor functions, 31 participants (17 female) were included. The gestational age at birth of the sample ranged from 34 to 42 weeks. At the time of data analysis, no participant was diagnosed with a neurological or neuromotor deficit, nor any neurodevelopmental impairment. All parents of the participating infants provided written informed consent to study participation and publication of depersonalized data and results. The study was approved by the Institutional Review Board (Ethics Commission) of the University Medical Center Göttingen (19/2019) and performed in accordance with the relevant guidelines and regulations.
@@ 2.1.2 Infant movements recordings
Standard laboratory recordings of infant movements at three timepoints (T1 – T3) were included in the current study. These were extracted from data of the umbrella project, assessments, among others, infants’ spontaneous movements in a standard laboratory-setting from the 4th to the 18th week of post-term age (PTA; corrected age; from here on, ages refer to the post-term age if not otherwise specified). For the current study, available laboratory recordings (n = 75) of the participants for the following timepoints are analyzed: T1: 28 $\pm$ 2 days, T2: 84 $\pm$ 2 days, and T3: 112 $\pm$ 2 days of PTA. As known, infants at 4 weeks (corresponding to T1) present a different spontaneous movement pattern than at 12 and 16 weeks (corresponding to T2 and T3; [55, 56]). With standard recordings from T1 to T3, we intended to cover the distinct age-specific spontaneous movement repertoires of the young infants [6].
@@ 2.1.4 Image selection and annotation
The 75 recordings were split into two parts of 50 and 25 which were annotated by one research assistant each, using JARVIS Annotation Tool 1.2 [57]. Additionally, 10% of the dataset, regardless of the split, were double-annotated for human error and individual joint labeling difficulty estimation.
@@ 2.2.1 Generic pose estimation
was one of the first pose estimation frameworks used for extraction of movements for analysis and classification of infant motor functions [30, 58, 31]. We included this framework as a baseline method to evaluate the performance gain of state-of-the-art models. We used the standard body 25 model available from https://github.com/CMU-Perceptual-Computing-Lab/openpose.
@@
This framework is not primarily intended for accurate pose estimation, but rather for fast inference on mobile devices. It was interesting for our comparison as Google advertises it as suitable for Yoga and fitness applications and included 25000 images of fitness exercises, so we expected it to deal well with different viewing angles and poses. We used version 0.9.2.1 as available through pip.
@@
The HR in HRNet stands for high resolution. The particular version used (HRNet-w48 available through MMPose [59], https://mmpose.readthedocs.io/en/latest/model_zoo/body_2d_keypoint.html#cid-hrnet-on-coco) has an input size of $512\times 512$ pixels. Because our videos are of sufficient resolution and adequate quality with low compression artefacts, this could benefit the pose estimation accuracy.
@@
This model is the current best performing framework for human pose estimation on the COCO test-dev dataset and uses a Vision Transformer architecture instead of the conventional CNN. We used the version available through mmpose [59] of the “huge” variety (https://mmpose.readthedocs.io/en/latest/model_zoo/body_2d_keypoint.html#topdown-heatmap-vitpose-on-coco).
@@ 2.2.2 Infant pose estimation
The AggPose model is based on the Transformer architecture, like ViTPose, and is specifically trained for infant pose detection on a proprietary dataset. In general, the images in the dataset also stem from recordings used for GMA, with infants in supine position, like in our dataset. We used the version published by the authors on github (https://github.com/PediaMedAI/AggPose).
@@
Soualmi et al. also retrained infant pose estimation networks on a proprietary dataset of GMA recordings. Compared to our dataset, the images contain clinical equipment (e.g., tubes, electrodes, cables), leading to more visual clutter in the scenes. The authors retrained multiple networks, of which we chose the HRNet48, because it is the same architecture as the one we used for the comparison of the generic pose estimation frameworks. We used the version published by the authors (https://drive.google.com/drive/folders/1SEuTqrNdz6ubRGwMaUazil0BVOqf2cYw?usp=sharing.)
@@ 2.2.3 Retraining
We retrained the best performing framework ViTPose (see Section 3.2) on our labeled dataset using 5-fold cross-validation. We manually split the dataset such that all images of each infant were contained in one fold only. This ensures the networks do not get better test results by overfitting on specific infants, which are correlated across different recording dates. Each of the five splits therefore contained 900 frames, made up from 450 different situations recorded from two different angles. Because of the different numbers of recordings done for each infant, four splits contained images from six different infants and one split contained images from seven different infants. For training, we randomly split 10% of the training set off as validation set, yielding a final training set size of 3240 frames in each fold. For each fold, we then retrained the model using the default training parameters for ViTPose from the MMPose framework [59]. Our graphics card did not have enough memory to retrain ViTPose-huge, so we retrained the smaller version ViTPose-large instead. We used the pretrained ViTPose model weights for initialization and employed a validation stop on PCK. We then kept the model with the highest PCK on the validation set for evaluation on the test set. The fold with the longest retraining took only 49 epochs until validation stop. Compared to the training of 210 epochs on the substantially larger COCO dataset that went into the base model, this is a relatively inexpensive retraining. All five models are published on Zenodo (https://doi.org/10.5281/zenodo.14833182).
The results presented for ”Retrained ViTPose” are the combined test set estimations of the five individual models, which in total yield the full dataset.
2.3 Quantification metrics
We used two main evaluation metrics to assess the performance of the pose estimation methods: the difference between predicted keypoint position and the human annotation, $d_{a}$, in pixels, and the percentage of correct keypoints (PCK).
Our dataset is not biased to any specific side (left or right) of the infants. Moreover, labeling left and right sides of hands, knees, etc., is not inherently different. Therefore, we merged the evaluation results for keypoints of the left and right side variety into one.
2.3.1 Pose estimation error
The difference between model prediction and the human annotation ($d_{a}$) is the Euclidean distance between the predicted and human label in pixels:
$d_{a}=\sqrt{(p_{x}-h_{x})^{2}+(p_{y}-h_{y})^{2}},$ (1)
where $p_{x/y}$ are the network predictions and $h_{x/y}$ the human labels for a keypoint.
2.3.2 Percentage of Correct Keypoints (PCK)
For PCK [60], all keypoint predictions are classified as either correct or incorrect, based on their distance to the label (as given above). The PCK score is then defined as the fraction of correct keypoints out of all keypoints. The distance threshold to be considered correct is defined as a percentage of the apparent torso length of the infant in the frame, measured as the distance between the left shoulder and hip. In this work we considered the PCK at 5%, 7.5% and 10% of the torso size, denoted by PCK@0.05, PCK@0.75 and PCK@0.1, respectively.
2.3.3 Relation to real-world distance
It is not possible to transform any of the quantification measures directly into real-world distances without using 3D triangulation, which was not done in this work, or assuming the points are on the same plane, which is untrue because the infants frequently lift their extremities. For better intuition, we can still provide an upper limit estimation (i.e., the real world distance can not be greater than this) to relate our evaluation metrics to the real-world distance in our recording setup. In our setting, the size of one pixel on the plane of the bed, always corresponds to a distance $\leq 0.8$ mm (in any camera view). Therefore, 80% of the pixel difference is an upper bound to the real-world distance in mm. Considering the mean torso length of 306 pixels, PCK@0.1 yields the percentage of points that were detected within approximately 2.5 cm of the ground truth.